OpenStudy (precal):

Jeremy has a box of 500 nails that weighs 1.35 kilograms. He uses 60 nails to build a birdhouse. How much do the nails in the birdhouse weigh?

OpenStudy (precal):

6th grade problem.......yuck

OpenStudy (precal):

I probably should convert the 1.35

OpenStudy (p0sitr0n):

500 nails = 1.35 kg ----------------- 60 nails = x x = 1.35kg * 60 nails / 500 nails = 0.162 kg Weight = m * g Weight = 0.162 * 9.8m/s^2 Weight = 1.59 N

OpenStudy (precal):

what does N stand for? Excuse my ignorance, we don't use metric in the US

OpenStudy (p0sitr0n):

newtons. wait, u need the mass or the weight ? Ok if its 6th grade, i assume ur answer is 0.162 kg

OpenStudy (precal):

yes kg Thanks :)